OpenStudy (anonymous):

How would you evaluate the integral of dx/(a^2+x^2) by first changing it into a more familiar integral using the substitution x=au?

terenzreignz (terenzreignz):

Let's do what the question suggests :) \[\Large \int \frac{dx}{a^2+x^2}\]

terenzreignz (terenzreignz):

It says let x = au right?

terenzreignz (terenzreignz):

Following that, must be dx?

OpenStudy (anonymous):

When I tried this myself I ended up with that substitution and du = (1/a)dx but when I plugged things in it wasn't turning out for me. I know it's supposed to end up as (1/a) arctan(x/a) but I have no idea where the arctan comes from.

terenzreignz (terenzreignz):

Well, to change it to 'something more familiar', it would seem you actually have to be familiar with it first :D Having said that, you know what the derivative of arctan is? \[\Large \frac{d}{dx}\tan^{-1}(x) = \color{red}?\]

OpenStudy (anonymous):

Yes, it's \[1/(1 + x^2) \]

terenzreignz (terenzreignz):

Oh... okay, great :D Now, after we let x = au it follows that dx = a du right?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Right, let's make that substitution then... \[\large x = \color{red}{au}\]\[\large dx = \color{blue}{a \ du}\] \[\Large \int \frac{\color{blue}{a \ du}}{a^2 + (\color{red}{au})^2}\]

terenzreignz (terenzreignz):

Catch me so far?

OpenStudy (anonymous):

yep I got that far

terenzreignz (terenzreignz):

Okay, at the denominator, we can 'expand', and then factor out: \[\Large \int \frac{a \ du }{a^2 + a^2u^t}\]

terenzreignz (terenzreignz):

\[\Large \int \frac{a \ du}{a^2(1+u^2)}\]

terenzreignz (terenzreignz):

Can you take it from here?

OpenStudy (anonymous):

I think so, let me look a bit.

terenzreignz (terenzreignz):

Remember that a is just a constant...

OpenStudy (anonymous):

Yeah I got it now, thank you so much!

terenzreignz (terenzreignz):

Okay, awesome ^_^